For an enterprise is given the following table, each worker earns a daily compensation of $80.


Number of Labour TP AP MP

0 0

1 8

2 14

3 18

4 25

5 30

6 32



(I) Calculate AP, MP, AVC, MC, and show your calculations.


(II) Represent in two different graphs the AP and MP(first graph) and AVC and MC (second graph).

(III) Compare MP and MC curves. Indicate similarities and differences.


(IV) Find at which point labor productivity starts to decline and explain why this happens?



1
Expert's answer
2022-02-24T11:42:17-0500

"AP=\frac{TP}{L}"

"MP=\frac{TP_2-TP_1}{L_2-L_1}"

"TC= L\u00d780"

"AVC=\frac{TC}{TP}"

"MC= \frac{TC_2-TC_1}{TP_2-TP_1}"